OpenStudy (anonymous):

In a game, a participant pays Rs. 20 to roll a dice. If he gets a prime number, he wins Rs. 30; if he gets a composite number, he loses his money; and if he gets neither prime nor composite, he gets Rs. 50. What is his average profit percentage?

OpenStudy (anonymous):

@queelius

OpenStudy (anonymous):

prime-2,3,5 so 1/2-->30

OpenStudy (anonymous):

composite 4,6 so 1/3 --> lose -20

OpenStudy (anonymous):

gets neither prime nor composite so 1 --1/6--->rs50

OpenStudy (anonymous):

30/2-20/38+50/6=16.66

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dg2 The integers between 1 and six may be sorted into these three sets. Prime: {2, 3, 5}, three in number Composite: {4, 6}, two in number Neither: {1}, one in number P(prime) = 3/6 = 1/2 P(composite) = 2/6 = 1/3 P(neither) = 1/6 This is the expectation of the profit. E[profit] = (1/2)(30 - 20) + (1/3)(-20) + (1/6)(50 - 20) = 5 - 20/3 + 5 = 10/3 The expectation of the profit is Rs. 10/3. For the expectation of the profit percentage that must be divided by the investment. (10/3)/20 = 1/6 = (100/6)% ≈ 16.7% Hope it helps